COFFEE AT $3.75 PER POUND IS BLENDED WITH COFEE AT $4.25 A POUND. HOW MUCH OF EACH MUST BE USED TO MAKE 50 POUNDS OF A MIXTURE AT $3.95 A POUND?
I HAVE 3.75X+4.25X=197.50
THAN ADD MY TWO X'S 8.00X=197.5
DIVIDE EACH BY 8.00 AND I GET 24.69 THIS ISNT RIGHT HUH???

Answer by stanbon(75887)

COFFEE AT $3.75 PER POUND IS BLENDED WITH COFEE AT $4.25 A POUND. HOW MUCH OF EACH MUST BE USED TO MAKE 50 POUNDS OF A MIXTURE AT $3.95 A POUND?
-----------------------
Value Equation: 3.75x + 4.25(50-x) = 3.95*50
375x + 425*50 - 425x = 395*50
-50x = -30*50
x = 30 lbs. (# of lbs of 3.75 coffee in the mixture)
50-x = 20 lbs. (# of lbs of 4.25 coffee in the mixture)
===================================================
Cheers,
Stan H.